# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v2

@@ 4.2 Sensitivities of poles
In this subsection, we consider sensitivities of poles $z_{j}$ of the reconstruction problem (1.1) that is solved with the method described in Subsection 4.1. First we consider the case of unstructured perturbations. Poles $z_{j_{n}}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$, are reconstructed as eigenvalues of the matrix pencil (4.12). Since Hankel matrices in the matrix pencil (4.12) have similar structure as Hankel matrices in the matrix pencil (2.6), we apply ideas from Subsection 3.1 without any changes and obtain for the sensitivities $\rho_{n}^{(-)}$, $n=1,\ldots,M_{1}$ of the poles $z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}$ in the $\mathbb{D}$ the formula for the exact computation
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
×$\times\left(|z_{j_{n}}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.14)
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{1}}$ and the upper bound
@@
$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left(\left|z_{j_{n}}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}z_{j_{k}}^{(-)}\right|\right).$
Now consider the case of the sensitivities $\rho_{n}^{(+)}$ of poles $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ located outside the unit circle $\mathbb{D}$. Inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of poles $z^{(+)}_{j_{n}}$ are reconstructed as eigenvalues of the matrix pencil (4.13). That means that we will compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of the poles $z^{(+)}_{j_{n}}$, but we will say that $\rho_{n}^{(+)}$ are also sensitivities of poles $z^{(+)}_{j_{n}}$.
To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results of the Subsection 3.1. Matrices $\mathbf{H}^{(+)}_{M_{2}}(0)$ and $\mathbf{H}^{(+)}_{M_{2}}(1)$ satisfy factorizations
@@
Thus left and right eigenvectors of the GEP of the Hankel matrix pencil (4.13) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{j_{n}}}$ are given by formulas
@@
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{2}}$ Taking into account (4.16) and (4.18), we get
@@
$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}.$ (4.19)
From (3.4) applying (4.16)–(4.19), we obtain the formula for the exact computation of sensitivities $\rho_{n}^{(+)}$ of the poles $z^{(+)}_{j_{n}}$ (the eigenvalues $\frac{1}{z^{(+)}_{j_{n}}}$ of the matrix pencil (4.13)) located outside the unit circle
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
×$\times\bigg{(}\frac{1}{|z_{j_{n}}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg{)},$ (4.20)
and the upper bound
$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{j_{n}}^{(+)}|}\,\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}+\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
Now since
$\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|,\ \ \ \big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|,$
we get for $n=1,\ldots,M_{2}$
@@
with
$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\left(\frac{1}{|z_{j_{n}}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|\right).$
Define further two vectors of unstructured sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$. Then using (4.15) and (4.21), we write the following upper bounds for the $\ell_{2}$-norm of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
@@
Using ideas similar as in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norm of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.23)
@@
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$. Naturally, arises the question how sharp are the obtained inequalities (4.224.24). In Example 5.3, we present the numerical computations for the exact values of the $\ell_{2}$- and $\ell_{1}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bound. Based on the results presented in Table 10, we can conclude that estimates (4.224.24) are quite sharp.
@@
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k,n}$
$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.25)
@@
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k,n}$
$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.26)
Formulas (4.25) and (4.26) are obtained using similar ideas from Subsection 3.2. In (4.26), we take into account (4.19).
Now we define a structured sensitivity of the pole $z^{(-)}_{j_{n}}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$. Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define a structured sensitivity of the pole $z^{(+)}_{j_{n}}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$. Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case. We define also two vectors
@@
are structured sensitivities for the poles $z_{j_{n}}^{(-)}$ and $z_{j_{n}}^{(+)}$ respectively.
Table 1: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.1

### v3

@@ 4.2 Sensitivities of poles
In this subsection, we present sensitivity analysis for poles $z_{j}$ of the rational function (1.1), reconstructed with the method described in Subsection 4.1. First, we consider the case of unstructured perturbations. Poles $z_{n}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$ are reconstructed as eigenvalues of the matrix pencil (4.12). Since Hankel matrices in (4.12) have structures similar to those of Hankel matrices in (2.6), we apply the ideas from Subsection 3.1 without any changes and obtain the formula for the exact computation of sensitivities $\rho_{n}^{(-)}$ with respect to unstructured perturbations of the poles $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$,
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{n}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
$\times\left(|z_{n}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.14)
where $\boldsymbol{e}_{n}$ is the $n$th canonical column vector in $\mathbb{R}^{M_{1}}$. Correspondingly, for the upper bound of $\rho_{n}^{(-)}$ we have
@@
$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{n}^{(-)}\right|}\left(\left|z_{n}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{k}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{k}^{(-)}z_{k}^{(-)}\right|\right).$
Now consider the case of sensitivities $\rho_{n}^{(+)}$ with respect to unstructured perturbations for the poles $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$ located outside $\mathbb{D}$. As was already mentioned, the inverse $\frac{1}{z^{(+)}_{n}}$ of the poles $z^{(+)}_{n}$ are reconstructed as eigenvalues of the matrix pencil (4.13). Thus, we compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{n}}$ of the poles $z^{(+)}_{n}$, but we say that $\rho_{n}^{(+)}$ are also sensitivities of $z^{(+)}_{n}$.
To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results from Subsection 3.1. Matrices $\mathbf{H}^{(+)}_{M_{2}}(0)$ and $\mathbf{H}^{(+)}_{M_{2}}(1)$ satisfy factorizations
@@
Thus, right and left eigenvectors of the GEP of the Hankel matrix pencil (4.13) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{n}}$ are given by formulas
@@
respectively, where $\boldsymbol{e}_{n}$ is the $n$th canonical column vector in $\mathbb{R}^{M_{2}}$ . Taking into account (4.16) and (4.18), we get
@@
$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{n}^{(+)}\,(z_{n}^{(+)})^{-2}.$ (4.19)
[…]
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{n}^{(+)}\,(z_{n}^{(+)})^{-2}|}\bigg\|\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg\|_{2}^{2}$
$\times\bigg(\frac{1}{|z_{n}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg),$ (4.20)
and the corresponding upper bound
$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{n}^{(+)}|}\,\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big\|_{2}+\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big\|_{2}}{|\gamma_{n}^{(+)}\,(z^{(+)}_{n})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
Since
$\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big\|_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-2}|,\ \ \ \big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big\|_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-3}|,$
the last inequality implies
@@
where
$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{n}^{(+)}\,(z^{(+)}_{n})^{-2}|}\left(\frac{1}{|z_{n}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-3}|\right).$
Define further two vectors of sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$. Then using (4.15) and (4.21), we write the following upper bounds for the $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
@@
Applying ideas similar to those in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norms of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2},$ (4.23)
@@
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$. Naturally, arises the question of how sharp are the inequalities (4.22)–(4.24). This question is discussed in Example 5.3.
@@
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}$:= $\frac{1}{\gamma_{n}^{(-)}}\bigg(\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{m-k,n}$
$-z_{n}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{m-k+1,n}\bigg)$ (4.25)
@@
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}$:= $-\frac{1}{\gamma_{n}^{(+)}\,z_{n}^{-2}}\bigg(\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{m-k,n}$
$-\frac{1}{z_{n}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{m-k+1,n}\bigg).$ (4.26)
Formulas (4.25) and (4.26) are obtained using ideas similar to those presented in Subsection 3.2. In (4.26) we take into account (4.19).
Now we define the sensitivity $\eta_{n,m}^{(-)}$ of the pole $z^{(-)}_{n}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$. Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define the sensitivity of the pole $z^{(+)}_{n}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$. By
$(\boldsymbol{\eta}^{(n)}_{2M_{1}})^{(-)}:=(\eta_{n,1}^{(-)},\ldots,\eta_{n,2M_{1}}^{(-)})^{T}\text{ and }(\boldsymbol{\eta}^{(n)}_{2M_{2}})^{(+)}:=(\eta_{n,1}^{(+)},\ldots,\eta_{n,2M_{2}}^{(+)})^{T}$ (4.27)
we denote the corresponding sensitivity vectors. Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case. We consider also two vectors
@@
are sensitivities of the poles $z_{n}^{(-)}$ and $z_{n}^{(+)}$, respectively, with respect to structured perturbations.
Table 1: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.1